A block slides down a ramp from rest. The block starts at 5.71 m off the ground and has a mass of 3.37 kg. Assuming no friction, what is the speed of the block at the bottom of the ramp?


1
Expert's answer
2021-06-24T16:14:52-0400

"mgh=mv^2/2\to v=\sqrt{2gh}=\sqrt{2\cdot9.81\cdot5.71}=10.58\ (m/s)" . Answer